# Multi-Rank Subspace Change-Point Detection for Monitoring Robotic Swarms

### v1

@@ Appendix A Proofs
$=\mbox{tr}\Big{[}(\sigma^{2}I_{k}+U\Lambda U^{T})\Big{(}\frac{1}{w}\Xi_{i}+u_{i}u_{i}^{T}\Big{)}\Big{]}+o\Big{(}\frac{1}{w}\Big{)}$
$=\sigma^{2}(1+\rho_{i})\Big{[}1+\frac{k-1}{w\rho_{i}^{2}}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{d}\frac{(1+\rho_{j})^{2}}{w(\rho_{i}-\rho_{j})^{2}}\Big{]}+o\Big{(}\frac{1}{w}\Big{)}.$
Therefore, we obtain
$\begin{split}\mathbb{E}_{0}Z_{t}&=\mathbb{E}_{0}\sum_{i=1}^{d}(\hat{u}_{i}^{T}x_{t})^{2}\\ &=\sigma^{2}\sum_{i=1}^{d}(1+\rho_{i})+\frac{\sigma^{2}}{w}\sum_{i=1}^{d}(1+\rho_{i})\Big{[}\frac{k-1}{\rho_{i}^{2}}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{d}\frac{(1+\rho_{j})^{2}}{(\rho_{i}-\rho_{j})^{2}}\Big{]}+o\big{(}\frac{1}{w}\big{)}.\end{split}$
@@ B.1 Comprehensive simulation results
Table A.1: Rank-1 dense spike: ARL = 5000, $\lambda=1$, $u=\frac{1}{\sqrt{k}}(1,1,\dots,1)^{T}$.
$k$ $\sigma^{2}=2$ $\sigma^{2}=1$ $\sigma^{2}=0.5$
CUSUM S-CUSUM EigChart CUSUM S-CUSUM EigChart CUSUM S-CUSUM EigChart
5 90.1 211.4 550.3 35.5 97.7 136.2 14.7 65.6 57.4
10 90.6 303.2 909.7 35.1 119.1 167.1 14.7 74.3 66.4
20 90.5 644.1 2194.4 35.5 136.4 203.5 14.7 80.1 72.9
Table A.2: Rank-1 sparse spike: ARL = 5000, $\lambda=1$, $u=e_{1}$.
$k$ $\sigma^{2}=2$ $\sigma^{2}=1$ $\sigma^{2}=0.5$
CUSUM S-CUSUM EigChart CUSUM S-CUSUM EigChart CUSUM S-CUSUM EigChart
5 90.9 230.6 561.0 35.2 102.4 141.2 14.8 69.5 59.9
10 90.3 411.5 911.4 35.6 112.8 161.2 14.9 70.3 62.1
20 90.4 788.7 2255.1 35.4 145.8 198.4 14.8 73.0 68.7
Table A.3: Rank-$d$ non-uniform spikes: ARL = 5000, $\Lambda=\mbox{Diag}(\text{linspace}(2,1,d))$.
@@
CUSUM S-CUSUM EigChart CUSUM S-CUSUM EigChart CUSUM S-CUSUM EigChart
5 2 28.6 90.5 324.7 11.6 63.9 67.2 5.3 56.0 38.9
5 3 20.8 76.9 256.6 8.4 61.0 63.8 3.9 54.2 35.5
10 2 28.5 109.4 581.1 11.6 65.5 76.5 5.3 56.7 43.2
10 3 20.8 90.9 403.0 8.3 59.9 71.2 3.7 54.5 40.4
20 2 28.5 153.0 1152.4 11.6 72.1 102.4 5.3 56.4 44.9
20 3 20.8 102.9 956.4 8.3 64.7 92.6 3.9 55.1 41.7

### v2

@@ Appendix A Proofs
$=\mbox{tr}\Big[(\sigma^{2}I_{k}+U\Lambda U^{T})\Big\{\frac{1}{w}\Xi_{i}+\Big(1-\frac{\mbox{tr}(\Xi_{i})}{w}\Big)u_{i}u_{i}^{T}\Big\}\Big]+o\Big(\frac{1}{w}\Big)$
$=\mbox{tr}\Big[(\sigma^{2}I_{k}+U\Lambda U^{T})\Big\{\frac{1+\rho_{i}}{w\rho_{i}^{2}}(I_{k}-UU^{T})+\frac{1}{w}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{d}\frac{(1+\rho_{i})(1+\rho_{j})}{(\rho_{i}-\rho_{j})^{2}}u_{j}u_{j}^{T}$
$\qquad+\Big(1-\frac{1+\rho_{i}}{w\rho_{i}^{2}}(k-d)-\frac{1}{w}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{d}\frac{(1+\rho_{i})(1+\rho_{j})}{(\rho_{i}-\rho_{j})^{2}}\Big)u_{i}u_{i}^{T}\Big\}\Big]+o\Big(\frac{1}{w}\Big)$
$=\sigma^{2}(1+\rho_{i})\Big\{1-\frac{k-d}{w\rho_{i}}-\frac{1}{w}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{d}\frac{1+\rho_{j}}{\rho_{i}-\rho_{j}}\Big\}+o\Big(\frac{1}{w}\Big).$
Therefore, summing over $i$ from $1$ to $d$, the double sum vanishes by antisymmetry, and we obtain the expected post-change increment by
$\begin{split}\mathbb{E}_{0}Z_{t}&=\sum_{i=1}^{d}\mathbb{E}_{0}(\hat{u}_{i}^{T}x_{t})^{2}\\ &=\sigma^{2}\sum_{i=1}^{d}(1+\rho_{i})\Big(1-\frac{k-d}{w\rho_{i}}\Big)-\frac{\sigma^{2}}{w}\sum_{i=1}^{d}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{d}\frac{(1+\rho_{i})(1+\rho_{j})}{\rho_{i}-\rho_{j}}+o\Big(\frac{1}{w}\Big)\\ &=\sigma^{2}\sum_{i=1}^{d}(1+\rho_{i})\Big(1-\frac{k-d}{w\rho_{i}}\Big)+o\Big(\frac{1}{w}\Big).\end{split}$
Proof of Proposition 4.5.
Suppose we are given the $\delta_{\infty}=\delta_{\infty}(\Delta)$ that satisfies $\mathbb{E}_{\infty}[e^{\delta_{\infty}(Z_{t}-\Delta)}]=1$. Then, by (tartakovsky2014sequential, Chapter 3), as we impose the ARL $\mathbb{E}_{\infty}[T^{\mathrm{sub}}]=\gamma\to\infty$, we have
$\mathbb{E}_{0}[T^{\mathrm{sub}}]=\frac{\log\gamma(1+o(1))}{\delta_{\infty}(\mathbb{E}_{0}Z_{t}-\Delta)}+w.$ (A.5)
Let $\tilde{\Sigma}=(\sigma^{-2}I_{k}-2\delta_{\infty}\hat{U}_{t}\hat{U}_{t}^{T})^{-1}$. Then, $\big\lvert\tilde{\Sigma}\big\rvert=\sigma^{2k}/(1-2\sigma^{2}\delta_{\infty})^{d}$ and thus
$1=\mathbb{E}_{\infty}\big[e^{\delta_{\infty}(Z_{t}-\Delta)}\big]$ $=e^{-\delta_{\infty}\Delta}\mathbb{E}_{\infty}\big[\mathbb{E}_{\infty}[e^{\delta_{\infty}Z_{t}}\big|\hat{U}_{t}]\big]$
$=e^{-\delta_{\infty}\Delta}\mathbb{E}_{\infty}\Bigg[\int_{\mathbb{R}^{k}}e^{\delta_{\infty}x_{t}^{T}(\hat{U}_{t}\hat{U}_{t}^{T})x_{t}}\frac{e^{-x_{t}^{T}x_{t}/(2\sigma^{2})}}{\sqrt{(2\pi)^{k}\sigma^{2k}}}dx_{t}\Bigg]$
$=e^{-\delta_{\infty}\Delta}\mathbb{E}_{\infty}\Bigg[\frac{1}{\sqrt{(2\pi)^{k}\sigma^{2k}}}\Bigg(\int_{\mathbb{R}^{k}}\left|2\pi\tilde{\Sigma}\right|^{-1/2}e^{-x_{t}^{T}\tilde{\Sigma}^{-1}x_{t}/2}dx_{t}\Bigg)\left|2\pi\tilde{\Sigma}\right|^{1/2}\Bigg]$
$=e^{-\delta_{\infty}\Delta}(1-2\sigma^{2}\delta_{\infty})^{-d/2}.$
Therefore, it holds that
$\Delta=-\frac{d}{2\delta_{\infty}}\log(1-2\sigma^{2}\delta_{\infty}).$ (A.6)
Using Lemma 4.3 and (A.5), we get
$\mathbb{E}_{0}[T^{\mathrm{sub}}]=\frac{\log\gamma(1+o(1))}{\sigma^{2}\delta_{\infty}A+\frac{d}{2}\log(1-2\sigma^{2}\delta_{\infty})}+w.$ (A.7)
The denominator of (A.7) is concave with respect to $\delta_{\infty}$, so the minimizer is given as
$\delta_{\infty}=\frac{1}{2\sigma^{2}}\Big(1-\frac{d}{A}\Big).$
Substituting this minimizer to (A.7) and (A.6), we get (15) and (13) respectively. Also, as we assume $w=o(\log\gamma)$, the threshold $b$ can be computed asymptotically as $b={\log\gamma(1+o(1))}/{\delta_{\infty}}.$ This results in (12). ∎
Proof of Theorem 4.7.
Recall (5) that the Kullback-Leibler divergence is given by
$\mathcal{I}_{0}\coloneqq\mathbb{E}_{0}\log\left(\frac{f_{0}}{f_{\infty}}\right)=\frac{1}{2}\sum_{i=1}^{d}[\rho_{i}-\log(1+\rho_{i})].$
Therefore, by the results in tartakovsky2014sequential, we have
$\mathbb{E}_{0}[T_{C}]=\frac{\log\gamma}{\mathcal{I}_{0}}+O(1)=\frac{2\log\gamma}{\sum_{i=1}^{d}[\rho_{i}-\log(1+\rho_{i})]}+O(1).$ (A.8)
Using the result from Proposition 4.5 and the optimal window size $w^{*}\asymp\sqrt{\log\gamma}$ from Corollary 4.6, the EDD of the MRS-C is
$\mathbb{E}_{0}[T^{\mathrm{sub}}]=\frac{2\log\gamma}{A(w^{*})-d(1+\log(\frac{A(w^{*})}{d}))}+w^{*}+O(1),$
and the denominator term approaches its limit
$D^{\infty}\coloneqq\lim_{\gamma\to\infty}\left[A(w^{*})-d\Big(1+\log\big(\frac{A(w^{*})}{d}\big.)\Big)\right]=\sum_{i=1}^{d}[\rho_{i}-\log(1+\bar{\rho})],$
with an error of order $O(1/\sqrt{\log\gamma})$. Therefore, the EDD for MRS-C is expanded as
$\mathbb{E}_{0}[T^{\mathrm{sub}}]=\frac{2\log\gamma}{\sum_{i=1}^{d}[\rho_{i}-\log(1+\bar{\rho})]}+O\big(\sqrt{\log\gamma}\big).$ (A.9)
Using the geometric series expansion $(1+x)^{-1}=1-x+O(x^{2})$, we get
$\frac{\mathbb{E}_{0}[T^{\mathrm{sub}}]}{\mathbb{E}_{0}[T_{C}]}$ $=\left(\frac{2\mathcal{I}_{0}}{D^{\infty}}+O\Big(\frac{1}{\sqrt{\log\gamma}}\Big)\right)\left(1-O\Big(\frac{1}{\log\gamma}\Big)\right)$
$=\mathcal{K}+O\Big(\frac{1}{\sqrt{\log\gamma}}\Big).$
Since $x\mapsto\log(1+x)$ is strictly concave, Jensen’s inequality states that
$\frac{1}{d}\sum_{i=1}^{d}\log(1+\rho_{i})\leq\log(1+\bar{\rho}).$
Therefore, $\mathcal{K}\geq 1$, with equality holding if and only if all $\rho_{i}$ are equal. ∎
@@ B.1 Comprehensive simulation results
Table A.1: Rank-1 dense spike: $\text{ARL}=5000$, $\lambda=1$, $u=(1,1,\dots,1)^{T}/\sqrt{k}$.
@@
CUSUM MRS-C LESC CUSUM MRS-C LESC CUSUM MRS-C LESC
5 1 90.1 ($0.84$) 211.4 ($3.36$) 550.3 ($7.53$) 35.5 ($0.27$) 97.7 ($1.98$) 136.2 ($1.45$) 14.7 ($0.13$) 65.6 ($0.27$) 57.4 ($0.28$)
10 1 90.6 ($1.23$) 303.2 ($7.42$) 909.7 ($30.04$) 35.1 ($0.30$) 119.1 ($1.92$) 167.1 ($1.16$) 14.7 ($0.21$) 74.3 ($0.71$) 66.4 ($0.41$)
20 1 90.5 ($0.90$) 644.1 ($28.73$) 2194.4 ($66.08$) 35.5 ($0.12$) 136.4 ($1.59$) 203.5 ($3.34$) 14.7 ($0.14$) 80.1 ($1.03$) 72.9 ($0.39$)
Table A.2: Rank-1 sparse spike: $\text{ARL}=5000$, $\lambda=1$, $u=e_{1}$.
$k$ $d$ $\sigma^{2}=2$ $\sigma^{2}=1$ $\sigma^{2}=0.5$
CUSUM MRS-C LESC CUSUM MRS-C LESC CUSUM MRS-C LESC
5 1 90.9 ($0.91$) 230.6 ($5.44$) 561.0 ($5.91$) 35.2 ($0.39$) 102.4 ($2.11$) 141.2 ($0.43$) 14.8 ($0.18$) 69.5 ($0.52$) 59.9 ($0.28$)
10 1 90.3 ($1.12$) 411.5 ($8.35$) 911.4 ($15.25$) 35.6 ($0.35$) 112.8 ($2.07$) 161.2 ($0.74$) 14.9 ($0.24$) 70.3 ($0.94$) 62.1 ($0.30$)
20 1 90.4 ($1.05$) 788.7 ($34.20$) 2255.1 ($52.51$) 35.4 ($0.21$) 145.8 ($2.40$) 198.4 ($1.98$) 14.8 ($0.19$) 73.0 ($1.38$) 68.7 ($0.26$)
Table A.3: Rank-$d$ non-uniform spikes: $\text{ARL}=5000$, $\Lambda=\operatorname{diag}(\operatorname{linspace}(2,1,d))$.
$k$ $d$ $\sigma^{2}=2$ $\sigma^{2}=1$ $\sigma^{2}=0.5$
CUSUM MRS-C LESC CUSUM MRS-C LESC CUSUM MRS-C LESC
5 2 28.6 ($0.15$) 90.5 ($0.57$) 324.7 ($4.65$) 11.6 ($0.11$) 63.9 ($0.17$) 67.2 ($0.46$) 5.3   ($0.02$) 56.0 ($0.05$) 38.9 ($0.18$)
5 3 20.8 ($0.11$) 76.9 ($0.32$) 256.6 ($2.51$) 8.4   ($0.04$) 61.0 ($0.10$) 63.8 ($1.25$) 3.9   ($0.00$) 54.2 ($0.05$) 35.5 ($0.25$)
10 2 28.5 ($0.23$) 109.4 ($1.10$) 581.1 ($2.60$) 11.6 ($0.03$) 65.5 ($0.27$) 76.5 ($1.95$) 5.3   ($0.03$) 56.7 ($0.08$) 43.2 ($0.20$)
10 3 20.8 ($0.09$) 90.9 ($0.53$) 403.0 ($0.35$) 8.3   ($0.04$) 59.9 ($0.05$) 71.2 ($2.11$) 3.7   ($0.01$) 54.5 ($0.07$) 40.4 ($0.14$)
20 2 28.5 ($0.23$) 153.0 ($1.57$) 1152.4 ($9.80$) 11.6 ($0.08$) 72.1 ($0.16$) 102.4 ($0.72$) 5.3   ($0.03$) 56.4 ($0.16$) 44.9 ($0.10$)
20 3 20.8 ($0.06$) 102.9 ($1.08$) 956.4 ($8.11$) 8.3   ($0.06$) 64.7 ($0.08$) 92.6 ($0.48$) 3.9   ($0.02$) 55.1 ($0.05$) 41.7 ($0.28$)